OpenStudy (anonymous):

What is the simplified form of the square root of the quantity 128 times n to the seventeenth power.? 8n8the square root of the quantity 2 times n. 4n8the square root of the quantity 8 times n. 2n8the square root of the quantity 8 times n. 8n8the square root of the quantity 4 times n.

OpenStudy (zzr0ck3r):

\[\sqrt{128n^{17}}=\sqrt{2^7*n^{16}n}=(2^6*2^1*n^{16}*n)^{\frac{1}{2}}=2^3n^8\sqrt{2n}=8n^8\sqrt{2n}\]

OpenStudy (anonymous):

thank you